A coffee merchant has coffee beans that sell for $9 per
pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?

Answer by checkley75(3666)

9x+12(100-x)=100*11.25
9x+1200-12x=1125
-3x=1125-1200
-3x=-75
x=-75/-3
x=25 pounds of $9 coffee.
then we need 75 pounds of $12 coffee.
proof
9*25+12*75=100*11.25
225+900=1125
1125=1125